A simple pendulum of length 30cm us suspended vertically. What is the period of the pendulum? What is the angular frequency of motion


1
Expert's answer
2021-12-12T16:39:56-0500

The period of the simple pendulum is given as follows:


"T = 2\pi \sqrt{\dfrac{L}{g}}"

where "L = 30cm = 0.3m" is its length, and "g = 9.8m\/s^2" is the gravitational acceleration.

Thus, obtain:


"T = 2\pi \sqrt{\dfrac{0.3m}{9.8m/s^2}} \approx 1.1s"

The angular frequency is:


"\omega =\sqrt{\dfrac{g}{L}} = \sqrt{\dfrac{9.8}{0.3}} \approx 5.7rad/s"

Answer. Period: 1.1s. angular frequency: 5.7 rad/s.